OpenStudy (anonymous):

A cyclist rides her bike at a speed of 10 meters per second. What is this speed in kilometers per hour? How many kilometers will the cyclist travel in 4 hours? Do not round your answers.

OpenStudy (greencat):

This is about conversion factors.

OpenStudy (anonymous):

yea

OpenStudy (greencat):

You need to know the conversion from meters to kilometers

OpenStudy (greencat):

Is work needed?

OpenStudy (anonymous):

1000 meters equals a kilometer

OpenStudy (greencat):

Like, can I shorten steps?

OpenStudy (greencat):

\[\frac{10m}{\sec}*\frac{1km}{1000m}\]

OpenStudy (greencat):

cancel out meters sign and multiply the fractions

OpenStudy (greencat):

What do you get after this?

OpenStudy (greencat):

you can also ccross out the 10 at the top

OpenStudy (greencat):

\[\frac{1km}{1000s}\]

OpenStudy (greencat):

But we're not done yet; you have to change the denominator to hours, which calss for a new conversion 60s=1min 60min=1hr 1hr=3600s

OpenStudy (greencat):

\[\frac{1km}{1000s}*\frac{3600s}{1hr}\] You can cancel out the seconds symbol and two zeroes: \[\frac{36km}{10hr}\]

OpenStudy (greencat):

^oops that seems to be the first answer

OpenStudy (greencat):

Note: I am solving so much because you are offline and I cannot properly guide you when you are not present.

OpenStudy (greencat):

The second part of the question asks for the distance traveled in 4hrs We can put what we know in the equation:\[Speed=\frac{Distance}{Time}\] We know the speed and the time so we can easily plug them in to solve for Distance.

OpenStudy (greencat):

\[\frac{36km}{10hr}=\frac{Distance}{4hr}\] Multiply both sides by 4hr \[\frac{36km*4hr}{10hr}=Distance\]

OpenStudy (greencat):

cancel out hr from top and bottom Also cancel out whatever you can in the fraction to get your answer Your answer to the second part should be in kilometers.

OpenStudy (greencat):

OS might mess up the order of my posts. If you need any more explanations, just ask me. Feel free to tag me in your future questions. ^_^

OpenStudy (greencat):

\[\Huge\color{green}{\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}\unicode{x1F63A}}\]